OpenStudy (anonymous):

Find the discontinuaties and zeros of (2x^2+5x-12)/(x+4)

OpenStudy (anonymous):

I got the discontinuity at (-4,-11) is that right? And I still need help with the zero(s).

OpenStudy (anonymous):

Okay, I got the zero of 3/2, is that correct?

OpenStudy (anonymous):

\[ \frac{2 x^2+5 x-12}{x+4}=\frac{(x+4) (2 x-3)}{x+4}=2 x-3 \]

OpenStudy (anonymous):

x =2/3 is correct

OpenStudy (anonymous):

x=-4 is the only removable discontinuity.

OpenStudy (anonymous):

How did you get the -11?

OpenStudy (anonymous):

That's where the point is on the graph, I wasn't stating another discontinuity.

OpenStudy (anonymous):

Ok. But -4 is a removable discontinuity.

OpenStudy (anonymous):

Alright, thank you.

OpenStudy (anonymous):

yw